OpenStudy (anonymous):

Difficult problem. I couldn't figure it out -- write in simplest radical form

OpenStudy (anonymous):

\[(\frac{ 25 }{ \sqrt[3]{20} })^{\frac{ 2 }{ 3 }}\]

OpenStudy (fibonaccichick666):

any ideas on how to start?

OpenStudy (anonymous):

Haha that's where I zoned out... do I have to bring the \[\sqrt[3]{20}\] to the top?

OpenStudy (fibonaccichick666):

no, distribute that 2/3 to start

OpenStudy (fibonaccichick666):

but first, change the cubed root to a fraction

OpenStudy (anonymous):

that would be \[20^{1/3}\]

OpenStudy (fibonaccichick666):

then?

OpenStudy (anonymous):

\[\frac{ 25^{2/3} }{ 20^{3/3} }\]

OpenStudy (fibonaccichick666):

stop, 20^(3/3)? check that

OpenStudy (anonymous):

Originally it was 20^1/3, I'm distributing the 2/3 power so shouldn't I add since it's the same base of 3

OpenStudy (fibonaccichick666):

do you add or multiply exponents?

OpenStudy (ranga):

\[ \Large (a^b)^c = a^{bc} \]

OpenStudy (ranga):

\[ \Large a^b \times a^c = a^{b+c} \]

OpenStudy (fibonaccichick666):

stahhpppp ranga

OpenStudy (anonymous):

haha so wait, that was correct?

OpenStudy (fibonaccichick666):

that's not necessary,mighty lion's knows that just made a mistake

OpenStudy (fibonaccichick666):

no mighty, your answer was not there yet. You have \[\frac{25^{2/3}}{[20^{1/3}]^{2/3}}\]

OpenStudy (fibonaccichick666):

correct? do you see?

OpenStudy (anonymous):

Ok I see. So now I raise 20^1/3 to the 2/3 power... geez this is complicated

OpenStudy (fibonaccichick666):

meh, not too, just apply the rules

OpenStudy (anonymous):

\[25^{\frac{ 2 }{ 3 }} \over 20^{\frac{ 2 }{ 9 }}\]

OpenStudy (fibonaccichick666):

good

OpenStudy (fibonaccichick666):

so now, change it back into roots

OpenStudy (anonymous):

Couldn't I just divide and get \[25^{\frac{ 4 }{ 9 }}\]

OpenStudy (anonymous):

Oh, I have to get it to roots because of simple radical form?

OpenStudy (fibonaccichick666):

no you can't divide

OpenStudy (fibonaccichick666):

no, because it makes it clearer for the next step

OpenStudy (anonymous):

So it would be \[\sqrt[3]{25^{2}} \over \sqrt[9]{20^{2}}\]

OpenStudy (fibonaccichick666):

good so now, compute 25^2 then find the cubed root if you can

OpenStudy (anonymous):

The cube root of 625 is \[5\sqrt[3]{5}\]

OpenStudy (anonymous):

Is \[5\sqrt[3]{5 } \over \sqrt[9]{20^2}\] the final answer?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (fibonaccichick666):

i bno, you have to simplify the bottom too

OpenStudy (anonymous):

\[\sqrt[9]{400}\]

OpenStudy (fibonaccichick666):

yea, now, can that be simplified any further?

OpenStudy (anonymous):

I don't think so.

OpenStudy (fibonaccichick666):

right, so we now want to try and make the top have a ninth root too

OpenStudy (fibonaccichick666):

so that we can actually do something interesting

OpenStudy (anonymous):

I don't know how to do that, never done a problem like this before

OpenStudy (fibonaccichick666):

the simplification you started with may have been too much lets go back to the \(^{3}\sqrt {625}\)

OpenStudy (fibonaccichick666):

so now for example, if I had 2 and I wanted it to be a cubed root, I would do this. \[^3\sqrt{2^3}\] you see how they undo each other and just leave you with two?

OpenStudy (fibonaccichick666):

so, same idea here, you want a ninth root

OpenStudy (fibonaccichick666):

but you have a 1/3

OpenStudy (anonymous):

Ah yes I see. So you add a number to both sides. So \[\sqrt[9]{625^6}\]

OpenStudy (fibonaccichick666):

do you follow?

OpenStudy (fibonaccichick666):

9/3=?

OpenStudy (anonymous):

3

OpenStudy (fibonaccichick666):

so it's \(625^?\)

OpenStudy (anonymous):

\[625^{3}\] right?

OpenStudy (fibonaccichick666):

yup, so now, remember how we distributed the 2/3 to start? we can now un-distribute(factor) the ninth root

OpenStudy (anonymous):

so \[625^{\frac{ 3 }{ 9 }}\]

OpenStudy (fibonaccichick666):

yea and now the whole fraction becomes?

OpenStudy (fibonaccichick666):

(once you get this I'l show you a faster way, I went really really roundabout)

OpenStudy (anonymous):

\[625^{\frac{ 3 }{ 9 }} \over \]\[\sqrt[9]{400}\]

OpenStudy (anonymous):

I can't take all night to figure this out though, I have like 30 problems after this

OpenStudy (fibonaccichick666):

yea, other way is faster jus undistribute the ninth root now then we can do the otehr way in 5 min

OpenStudy (fibonaccichick666):

it's the same principle which is what is important

OpenStudy (anonymous):

so what I just wrote is the correct answer?

OpenStudy (fibonaccichick666):

no,not yet

OpenStudy (fibonaccichick666):

Just restart, it takes like 2 min \[(\frac{ 25 }{ \sqrt[3]{20} })^{\frac{ 2 }{ 3 }}\] From the start, remember my example with 2 and cubed root of 2 cubed? Try and apply that to start

OpenStudy (fibonaccichick666):

(to the 25)

OpenStudy (anonymous):

\[\sqrt[3]{25^2}\]

OpenStudy (fibonaccichick666):

squared or cubed on the 25?

OpenStudy (anonymous):

Oh yeah, the other way around.

OpenStudy (fibonaccichick666):

so write the whole problem out